# Compass: A Comprehensive Tool for Accurate and Efficient Molecular Docking in Inference and Fine-Tuning

### v1

@@ Abstract
While there has been discussion about noise levels in molecular docking datasets such as PDBBind, a thorough analysis of their physical/chemical and bioactivity noise characteristics is still lacking. PoseCheck addresses this issue by examining molecular strain energy, molecular-protein clashes, and interactions, but it is primarily created for $de$ $novo$ drug design. Another important metric in molecular docking, Binding Affinity Energy, is better assessed by the new empirical score function, AA-Score, which has demonstrated improved performance over existing methods. To tackle these challenges, we propose the COMPASS method, which integrates the PoseCheck and AA-Score modules. This approach evaluates dataset noise levels and the physical/chemical and bioactivity feasibility of docked molecules. Our analysis of the PDBBind dataset using COMPASS reveals significant noise in the ground truth data. Additionally, we incorporate COMPASS with the state-of-the-art molecular docking method, DiffDock, in inference mode to achieve efficient and accurate assessments of docked ligands. Finally, we propose a new paradigm to enhance model performance for molecular docking through fine-tuning and discuss the potential benefits of this approach. The source code is available publicly at https://github.com/BIMSBbioinfo/Compass.
@@ 1 Introduction
Molecular docking is crucial in drug discovery as it helps determine the feasibility of potential drug candidates. Traditional docking methods are slow  (Trott & Olson, 2010; Halgren et al., 2004), leading to the development of deep learning (DL)-based methods  (Stärk et al., 2022; Lu et al., 2022). These newer approaches have shown impressive accuracy and significantly faster performance  (Corso et al., 2022; 2024; Lu et al., 2024).
Deep learning based methods evaluate their accuracy based on numerical metrics, such as Root Mean Square Distance (RMSD)  (Meli & Biggin, 2020); such metrics, however, do not accurately capture physico-chemical properties which characterize strong molecular binding. PoseCheck  (Harris et al., 2023) was created for Structure-Based Drug Design (SBDD)  (Blundell, 1996; Ferreira et al., 2015; Anderson, 2003). PoseCheck analyzes molecular stability through strain energy, finds interactions of complex, and counts steric clashes between the protein and ligand. Small molecule binding affinity energy is another crucial metric for evaluation of bioactivity of protein-ligand interactions. Autodock-Vina  (Trott & Olson, 2010) is a commonly used method to calculate the binding affinity energy. Recently, however, AA-Score  (Pan et al., 2022) - an empirical scoring function, has been shown as more accurate than Autodock-Vina. Additionally, AA-Score provides the structure of the docked ligand in the protein pocket region in 3D (in pdb file format), making it more effective for visual inspection of the docking results.
Additionally,  Corso et al. (2024) demonstrated that fine-tuning the model can boost performance. We believe that instead of training the model solely on physical features, allowing the model to learn the structural properties of protein-ligand interactions and then fine-tuning it with physical/chemical and bioactivity (PCB) constraints can make the model more efficient and better suited to real-world scenarios.
To tackle the challenges in DL-Based Molecular Docking, such as identifying ligand strain energy, counting protein-ligand clashes, finding interaction types, finding the best score function for binding affinity energy, and enhancing performance with a PCB-based approach, we propose COMPASS, a comprehensive tool with two modes:
  • Inference Mode: Provides comprehensive PCB-based analysis during DL-Based Molecular Docking with integrated COMPASS modules.
@@
Our analysis of the PDBBind  (Liu et al., 2017) dataset, commonly used in molecular docking, revealed significant PCB noise/violations. Also, secondly, we give users opportunity to analysis their docking with real assessments in inference run and finally, with these PCB-based assessments, we can improve model with fine-tuning with new score called Compass Score.
In summary, our study offers the following key contributions:
  • A PCB violation analysis of why DL methods do not achieve the desired performance, highlighting the noisy nature of the training dataset.
  • The development of a tool to PCB violation analyze docking results obtained through DL-based molecular docking in inference mode.
  • A new strategy on fine-tuning method for DL-Based Molecular Docking models with a physics-based approach that we called as Compass Score as penalizer.
Figure 1: (A): Inference Mode: Provides comprehensive PCB-based analysis during DL-Based Molecular Docking using integrated COMPASS modules during Inference. (B): Introduces a new fine-tuning method for DL-Based Molecular Docking with a PCB-based approach, utilizing the Integrated Compass Score to calculate for Binding Affinity (BA), Strain Energy (SE), and Number of Clashes (NC) using the LAN-MSE function. COMPASS: Consists of two modules: AA-Score, an empirical scoring function for calculating binding affinity energy and providing a 3D protein pocket, and PoseCheck for evaluating steric clashes, protein-ligand interactions, and ligand strain energy.
@@ 2 Methods
The proposed COMPASS method enhances molecular docking by integrating DiffDock a deep learning-based docking, PoseCheck for evaluating PCB properties like strain energy, protein-ligand interactions and steric clashes, and AA-Score for calculating binding affinity energy and analyzing protein pocket structures. By combining these modules, COMPASS provides a comprehensive and accurate assessment of molecular docking, addressing limitations of existing methods. Additionally, a new fine-tuning paradigm to improve the performance not just with ligand’s structure but also with their PCB properties. The overall structure of COMPASS is illustrated in Figure 1.
@@ 2.1 Motivation
Generally, methods are evaluated using the RMSD, with values below 2 Ångstroms considered nearly accurate (Alhossary et al., 2015; Hassan et al., 2017; McNutt et al., 2021). However, an analysis of physical violations using COMPASS, as shown in Figure 2, reveals that RMSD may not be a reliable metric for assessing the chemical and physical properties of docked molecules. Despite reasonable Binding Affinity and RMSD values in the exactly the same protein pocket, high Strain Energy and Steric Clash compared to the Ground Truth of PDB ID 1a46 suggest limitations. Nonetheless, RMSD remains useful for learning the geometry of molecules for docking due to its ease and speed of calculation during the training.
Also, when we have new protein-ligand pairs, we don’t always have Ground Truth Data. DiffDock Confidence Score  (Corso et al., 2022) tries to rank the conformations but what we discussed in RMSD also applies to Confidence Score.
To evaluate protein-ligand molecular docking correctly, we need a new tool that gives the molecule’s Binding Affinity, Strain Energy, Steric Clash with Protein and interaction types. To do this, we propose COMPASS that utilize PoseCheck and AA-Score modules.
Despite the advanced nature of deep learning (DL)-based methods, their performance improvements are quite limited, partly due to the noisiness of the PDBBind dataset used for training, which has not been adequately discussed or analyzed. Examination of approximately 14,000 protein-ligand pairs in the PDBBind Dataset using COMPASS reveals the presence of PCB violations within the dataset.
Finally,  (Corso et al., 2024) discuss and show how fine-tuning the model can enhance its performance. To improve performance, we introduce a new regularizer called Compass Score, which penalizes the loss function during fine-tuning based on PCB violations identified by our COMPASS tool.
2.2 COMPASS
As mentioned above, COMPASS consists of three modules: PoseCheck for explaining the physical interaction properties of ligand-protein; AAScore for determining binding affinity and identifying the docked pocket region; and DiffDock, a SoTA DL-based method integrated for accurate docking. All source codes are available  [^1] https://github.com/BIMSBbioinfo/Compass.
@@ 2.2.1 DiffDock
DiffDock  (Corso et al., 2022; 2024) approached molecular docking as a problem of learning a distribution of ligand poses based on the protein structure, utilizing a diffusion generative model (DGM) developed over this defined space. While  (De Bortoli et al., 2022) initially described DGMs on submanifolds by projecting a diffusion from ambient space onto the submanifold, they refined this by mapping to a more manageable manifold where the diffusion kernel is directly sampled, thereby enhancing the efficiency of training the DGM on this new manifold.
Ligand movements are mapped in molecular docking to mathematical groups, associating translations with the 3D translation group $T(3)$, rigid rotations with the 3D rotation group $SO(3)$, and changes in torsion angles with multiple copies of the 2D rotation group $SO(2)$. These movements are formalized by defining how each group acts on a ligand pose $c\in\mathbb{R}^{3n}$, where translations are straightforward additions to atom positions, and rotations are conducted around the ligand’s center of mass. Furthermore, modifications to torsion angles are defined to minimally perturb the structure in an RMSD sense, and can be specified separately from translations and rotations to ensure changes are disentangled.
Overall, DiffDock is comprised of two primary models: the Score Model and the Confidence Model.
@@ 2.2.2 PoseCheck
PoseCheck  (Harris et al., 2023) is one of the modules integrated into COMPASS to analyze the interactions of docked molecules by DiffDock. These analyses include:
@@
Steric Clashes: Steric clashes occur when two neutral atoms are closer than the sum of their van der Waals radii, indicating an energetically unfavorable and physically implausible interaction (Ramachandran et al., 2011; Buonfiglio et al., 2015). Such clashes suggest that the current conformation of the ligand within the protein is suboptimal, highlighting potential inadequacies in pose design or a fundamental mismatch in molecular topology. In structure-based drug design (SBDD), the total number of these clashes is considered a crucial performance metric, with a clash defined as the pairwise distance between protein and ligand atoms falling below their combined van der Waals radii by a tolerance of 0.5 Å  (Harris et al., 2023).
@@ 2.2.3 AA-Score
A new empirical scoring function developed by  (Pan et al., 2022), named AA-Score, which includes several amino acid-specific interaction components such as hydrogen bonds, electrostatic, and van der Waals interactions, illustrated in Equation 1:
$\sum_{i}\Delta G_{\text{vdW--main}}^{AA}$
where all components detailed in Appendix B.1. This scoring function differentiates interactions with main-chain and side-chain atoms across these types and incorporates additional energy components that are not amino acid-specific, including hydrophobic contacts, $\pi$-$\pi$ stacking, $\pi$-cation interactions, metal–ligand interactions, and a conformational entropy penalty. Each of these energy components is thoroughly described to provide a comprehensive understanding of their contributions to the scoring function.
2.3 Regalurized Loss with Compass Constant for Finetuning
In the latest version of DiffDock (Corso et al., 2024), it has been showned that fine-tuning enhances the performance of the model, even with small datasets. However, the fine-tuning still follows a structural framework, and PCB prior knowledges are overlooked.
In this context, using the COMPASS, which includes modules PoseCheck and AA-Score to calculate Binding Affinity, Strain Energy, and Number of Clash provides valuable information about PCB violations, making it a suitable basis for model enhancement. To better compare these against ground-truth data, we propose replacing traditional regression methods with a new scoring method called Log Absolute Normalized - Mean Square Error (LAN-MSE), which more effectively normalizes outliers (see Section 2.3.1).
@@ 2.3.1 Log Absolute Normalized - Mean Square Error (LAN-MSE)
The LAN-MSE score function applies a logarithmic transformation to both predicted and true values before computing the mean squared error (MSE). This logarithmic approach effectively reduces the impact of outliers, enhancing model robustness by prioritizing relative rather than absolute differences. This characteristic is particularly advantageous in scenarios where output values span large scales.
@@
where:
  • $\hat{y}_{i}$ represents the predicted value for the $i$-th data point.
  • $y_{i}$ is the true value for the $i$-th data point.
  • $\text{buffer}_{\text{true},i}$ is set to 1.1 if $|y_{i}|<1$, otherwise, it is 1.0.
  • $\text{buffer}_{\text{pred},i}$ is similarly set based on the magnitude of $\hat{y}_{i}$.
  • $\epsilon$ is a small constant (e.g., $10^{-5}$) to prevent division by zero in the denominator.
This detailed definition underscores how the application of buffers and normalization constants in the loss function ensures nuanced evaluation of prediction accuracy across various data magnitudes.(See Details for further Analysis in Appendix C)

### v2

# CompassDock [Uncaptioned image]: Comprehensive Accurate Assessment Approach for Deep Learning-Based Molecular Docking

@@ Abstract
Datasets used for molecular docking, such as PDBBind, contain technical variability - they are noisy. Although the origins of the noise have been discussed, a comprehensive analysis of the physical, chemical, and bioactivity characteristics of the datasets is still lacking. To address this gap, we introduce the Comprehensive Accurate Assessment (Compass). Compass integrates two key components: PoseCheck, which examines ligand strain energy, protein-ligand steric clashes, and interactions, and AA-Score, a new empirical scoring function for calculating binding affinity energy. Together, these form a unified workflow that assesses both the physical/chemical properties and bioactivity favorability of ligands and protein-ligand interactions. Our analysis of the PDBBind dataset using Compass reveals substantial noise in the ground truth data. Additionally, we propose CompassDock, which incorporates the Compass module with DiffDock, the state-of-the-art deep learning-based molecular docking method, to enable accurate assessment of docked ligands during inference. Finally, we present a new paradigm for enhancing molecular docking model performance by fine-tuning with Compass Scores, which encompass binding affinity energy, strain energy, and the number of steric clashes identified by Compass. Our results show that, while fine-tuning without Compass improves the percentage of docked poses with RMSD $<$ 2Å, it leads to a decrease in physical/chemical and bioactivity favorability. In contrast, fine-tuning with Compass shows a limited improvement in RMSD $<$ 2Å but enhances the physical/chemical and bioactivity favorability of the ligand conformation. The source code is available publicly at https://github.com/BIMSBbioinfo/CompassDock.
@@ 1 Introduction
Molecular docking is crucial in drug discovery as it helps determine the feasibility of potential drug candidates. Traditional docking methods are slow  (Trott & Olson, 2010; Halgren et al., 2004), leading to the development of deep learning (DL)-based methods  (Stärk et al., 2022; Lu et al., 2022). These newer approaches have shown impressive accuracy and significantly faster performance, decreasing the run time from hours/days to seconds/minutes per protein-ligand pair  (Corso et al., 2022; 2024; Lu et al., 2024).
Deep learning-based methods typically assess their accuracy using numerical metrics such as Root Mean Square Distance (RMSD)  (Meli & Biggin, 2020). Although RMSD measures the overall deviation of individual molecules from the ground truth, it is a purely distance-based metric and does not consider the physico-chemical interactions or bioactivity properties between the ligand and the protein at the binding site. As a result, metrics like RMSD fail to capture critical factors that determine strong molecular binding, such as the favorability of these interactions. PoseBuster  (buttenschoen2024posebusters) demonstrated that even when structures have RMSD $<$ 2Å, it is common to find ligands with unfavorable physico-chemical properties.
PoseCheck  (Harris et al., 2023) was developed for Structure-Based Drug Design (SBDD)  (Blundell, 1996; Ferreira et al., 2015; Anderson, 2003) to analyze molecular stability by calculating strain energy, identifying interactions within the complex, and counting steric clashes between the protein and ligand. Another critical metric for assessing protein-ligand interactions is binding affinity energy, which reflects bioactivity. Autodock-Vina  (Trott & Olson, 2010) is a widely used method for calculating binding affinity energy. However, the recently introduced AA-Score  (Pan et al., 2022), an empirical scoring function, has been shown to be more accurate than Autodock-Vina. Additionally, AA-Score provides the 3D structure of the docked ligand within the protein pocket (pdb file), allowing for easier visual inspection of docking results.
Additionally,  (Corso et al., 2024) demonstrated that model fine-tuning boosts performance, yielding an increased percentage of structures’ with an RMSD score of less than 2Å. While fine-tuning without Physico-Chemical and Bioactivity (PCB) features increases the percentage of docked poses with RMSD $<$ 2Å, we demonstrate that this approach results in degraded physico-chemical interactions and bioactivity. On the other hand, fine-tuning with PCB results in only a slight improvement in RMSD $<$ 2Å but enhances the ligands’ physico-chemical and bioactivity favorability.
To tackle the challenges in DL-based molecular docking, such as identifying ligand strain energy, counting protein-ligand steric clashes, finding interaction types, finding the best score function for binding affinity energy, and enhancing performance with a PCB-based approach, we propose CompassDock with two modes (Figure 1):
  • Inference Mode: Provides comprehensive PCB quality analysis during DL-based molecular docking with integrated Comprehensive Accurate Assessment (Compass) module.
@@
Figure 1: (A) Inference Mode: Provides comprehensive PCB feature quality assessment during DL-Based Molecular Docking Inference runtime using integrated Compass workflow. (B) Fine-Tuning Mode: Introduces a new fine-tuning method for DL-Based Molecular Docking with a PCB-based approach, utilizing the Integrated Compass Score.
Our analysis of the PDBBind  (Liu et al., 2017) dataset, commonly used in molecular docking, revealed significant PCB noise/unfavorability. Secondly, we give users the opportunity to analyze their docking results with real assessments in an inference run. Finally, with these PCB-based assessments, we can improve the model by fine-tuning it with a new score function called Compass Score.
In summary, our CompassDock offers the following key contributions:
  • A unified module, named Compass, was developed to determine PCB features of docked molecular conformations sampled by DL-based molecular docking during inference mode.
  • A PCB quality/favorability analysis of why DL methods do not achieve the desired performance, highlighting the noisy nature of the training dataset (i.e. PDBBind).
  • Proposing a novel loss function called Log Absolute Normalized - Mean Square Error (LAN-MSE) which prevents exploding loss values and effectively reduces the impact of outliers thereby enhancing model robustness.
  • A new objective function, called Compass Score, which incorporates binding affinity energy, strain energy, and steric clashes into LAN-MSE.
  • A new paradigm introduced to enhance molecular docking model performance in PCB quality by fine-tuning DL-based docking models using Compass Score.
@@ 2 Methods
The proposed CompassDock method not only enables accurate and realistic PCB quality assessments in DL-based molecular docking during inference, such as with DiffDock but also enhances the model’s ability to capture favorable PCB features by incorporating these properties into the fine-tuning process using a novel score (Compass Score) and loss function (LAN-MSE). The overall structure of CompassDock is illustrated in Figure 1.
@@ 2.1 Motivation
Generally, the performance of the docking methods is evaluated using the RMSD metric, where RMSD values below 2 Ångstroms(Å) are considered nearly accurate (Alhossary et al., 2015; Hassan et al., 2017; McNutt et al., 2021). However, PoseBuster  (buttenschoen2024posebusters) demonstrated that even in structures with RMSD $<$ 2Å, ligands can still display unfavorable physico-chemical properties. We further illustrated this by incorporating a bioactivity score during CompassDock inference, as shown in Figure 2, in order to highlight that the RMSD metric does not necessarily lead to physico-chemically and bioactively favorable conformation structures. Despite an RMSD of 1.23 within the same protein pocket region, the sampled molecular conformation exhibits less favorable binding affinity energy, along with higher strain energy and steric clashes compared to the ground truth of PDB ID 1a46.
Also, when we have new protein-ligand pairs, we don’t always have Ground Truth Data. DiffDock Confidence Score  (Corso et al., 2022) is used to rank the various conformations, but what we discussed in RMSD also applies to the Confidence Score.
To accurately evaluate protein-ligand molecular docking, a unified workflow integrated with DL-based docking methods, such as DiffDock, is needed to provide information on binding affinity, strain energy, steric clashes, and interaction types. To achieve this, we propose CompassDock, which utilizes the Compass module, combining PoseCheck and AA-Score, without the need for complex package management.
Despite advancements in deep learning (DL)-based molecular docking methods, their performance improvements remain limited, partly due to the noise present in the PDBBind dataset used for training, which has not been thoroughly discussed or analyzed. An examination of approximately 14,000 protein-ligand pairs in the PDBBind dataset using Compass reveals the presence of unfavorable PCB values, as shown in Figure 3.
Finally,  (Corso et al., 2024) showed that fine-tuning the model can improve the percentage of structures with RMSD $<$ 2Å. However, while fine-tuning without incorporating PCB features increases the proportion of docked poses achieving RMSD $<$ 2Å, our results demonstrated a decrease in the favorability of PCB features. To improve favorable PCB features, we introduce a new regularizer called Compass Score, which penalizes the loss function during fine-tuning based on PCB features.
2.2 CompassDock
As shown in Figure 1, CompassDock operates in two modes: Inference Mode (Section 2.3) and Fine-Tuning Mode(Section 2.4). In both modes, the core component is a unified module called Compass, which integrates with DL-based molecular docking methods, such as DiffDock, to assess how well molecular conformations exhibit favorable PCB features, including ligand strain energy, binding affinity energy, and the number of steric clashes between protein-ligand pairs. The Compass workflow/module consists of two key elements: PoseCheck and AA-Score. All source codes are available  [^1] https://github.com/BIMSBbioinfo/CompassDock.
@@ 2.2.1 DiffDock
DiffDock  (Corso et al., 2022; 2024) approaches molecular docking as a problem of learning a distribution of ligand poses based on the protein structure, utilizing a diffusion generative model (DGM) developed over this defined space. While  (De Bortoli et al., 2022) initially described DGMs on submanifolds by projecting a diffusion from ambient space onto the submanifold, they refined this by mapping to a more manageable manifold where the diffusion kernel is directly sampled, thereby enhancing the efficiency of training the DGM on this new manifold.
Ligand movements are mapped in molecular docking to mathematical groups, associating translations with the 3D translation group $T(3)$, rigid rotations with the 3D rotation group $SO(3)$, and changes in torsion angles with multiple copies of the 2D rotation group $SO(2)$. These movements are formalized by defining how each group acts on a ligand pose $c\in\mathbb{R}^{3n}$, where translations are straightforward additions to atom positions, and rotations are conducted around the ligand’s center of mass. Furthermore, modifications to torsion angles are defined to minimally perturb the structure in an RMSD sense and can be specified separately from translations and rotations to ensure changes are disentangled.
Overall, DiffDock consists of two primary models: the Score Model and the Confidence Model.
@@ 2.2.2 PoseCheck
PoseCheck  (Harris et al., 2023) is developed to assess molecular stability by measuring strain energy, identifying interactions within the complex, and counting steric clashes between the protein and ligand.
@@
Steric Clashes: Steric clashes occur when two neutral atoms are closer than the sum of their van der Waals radii, indicating an energetically unfavorable and physically implausible interaction (Ramachandran et al., 2011; Buonfiglio et al., 2015). Such clashes suggest that the current conformation of the ligand within the protein is suboptimal, highlighting potential inadequacies in pose design or a fundamental mismatch in molecular topology. In SBDD, the total number of these clashes is considered a crucial performance metric, with a clash defined as the pairwise distance between protein and ligand atoms falling below their combined van der Waals radii by a tolerance of 0.5 Å  (Harris et al., 2023).
@@ 2.2.3 AA-Score
A new empirical scoring function by  (Pan et al., 2022), named AA-Score, which includes several amino acid-specific interaction components such as hydrogen bonds, electrostatic, and van der Waals interactions (Appendix B.1). This scoring function differentiates interactions with main-chain and side-chain atoms across these types and incorporates additional energy components that are not amino acid-specific, including hydrophobic contacts, $\pi$-$\pi$ stacking, $\pi$-cation interactions, metal-ligand interactions, and a conformational entropy penalty. Each of these energy components is thoroughly described to provide a comprehensive understanding of their contributions to the scoring function.
2.3 Inference Mode
In inference mode, the protein (provided as either a sequence or 3D structure) and the ligand (given as a SMILES string or 3D structure) are input into DiffDock to generate 3D ligand conformations. Compass then analyzes the protein and ligand 3D structures to evaluate binding affinity, strain energy, the number of steric clashes, and protein interaction types. Additionally, we developed a recursive redocking method to further optimize the molecular conformation during inference, making the PCB features’ quality more favorable. This process is described by the following Equation 1:
$f(L_{i},P,n)=\begin{cases}L_{i}&\text{if the conformation is favorable}\\ \text{stop}&\text{if the conformation is unfavorable}\\ \text{stop}&\text{if }n\geq N_{\text{max}}\\ f(L_{i+1},P,n+1)&\text{otherwise (further refinement is needed)}\end{cases}$ (1)
where $L_{i}$ is current ligand conformation at step $i$; $P$ is the protein structure; $f$: the recursive function that refines ligand conformations based on PCB features; $n$ is the current iteration step and $N_{\text{max}}$ is the maximum number of iterations allowed.
This function iterates through docking steps, checking each conformation against the thresholds. It halts the recursion when the optimal conditions are met, when the scores are too low, or when the maximum iteration limit $N_{\text{max}}$ is reached. If none of these conditions are satisfied, it continues to the next conformation $L_{i+1}$, refining the results until a favorable state is achieved or the iteration limit is reached.
2.4 Fine-Tuning Mode
2.4.1 Loss Regularizer with Compass Score for Fine-Tuning
Fine-tuning of the DL-based docking method focuses on reducing the overall structural deviation of individual molecules from the ground truth but tends to overlook important PCB features’ quality of ligands and protein-ligand interactions. To address this, we used Compass to fine-tune the model with the goal of improving the quality of these PCB features. To more effectively compare these results with ground-truth data, we propose replacing traditional regression methods with a new loss function, Log Absolute Normalized - Mean Square Error (LAN-MSE), which better handles the normalization of outliers (Section 2.4.2).
@@ 2.4.2 Log Absolute Normalized - Mean Square Error (LAN-MSE)
The LAN-MSE loss function applies a logarithmic transformation to both predicted and true values before computing the mean squared error (MSE). This logarithmic approach effectively reduces the impact of outliers, enhancing model robustness by prioritizing relative rather than absolute differences. This characteristic is particularly advantageous in scenarios where output values span large scales.
@@
where $\hat{y}_{i}$ represents the predicted value for the $i$-th data point; $y_{i}$ is the true value for the $i$-th data point; $\text{buffer}_{\text{true},i}$ is set to 1.1 if $|y_{i}|<1$, otherwise, it is 1.0; $\text{buffer}_{\text{pred},i}$ is similarly set based on the magnitude of $\hat{y}_{i}$; $\epsilon$ is a small constant (e.g., $10^{-5}$) to prevent division by zero in the denominator.
This detailed definition underscores how the application of buffers and normalization constants in the loss function ensures a nuanced evaluation of prediction accuracy across various data magnitudes. (Details for further Analysis in Appendix C)